The velocity of an 11-kg box changes from 5.1 m/s to 15 m/s. How much total work is done on the box?
1
Expert's answer
2011-02-25T06:52:39-0500
The work would be equal to change of kinetic energy of the object:
A = mv22/2 - mv12/2 = m/2 (v22-v12) = 11/2 (152-5.12) = 1094.445 J.